OpenStudy (anonymous):

3-2{7x-[-2(3x-1)+3(5x-3^2)]} I feel like i got it buttttt get confused plzz hpl

Vocaloid (vocaloid):

start from the innermost parentheses/brackets and work your way outside let's start with -2(3x-1) = ?

OpenStudy (anonymous):

+6x + 2?

Vocaloid (vocaloid):

almost, check the sign on 6x

OpenStudy (anonymous):

i get confused with the negatives and positives..

OpenStudy (anonymous):

oh -6x

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

Can i show you how i wrote it?

Vocaloid (vocaloid):

negative*positive = negative

Vocaloid (vocaloid):

sure

OpenStudy (anonymous):

3-2 {7x-[-6x+2) + (15x-27)] 3+ - 14x + 12x -4 + -30x +54 i think im doing something wrong.. i dont know

OpenStudy (freckles):

you did all of that in one step? if so I think you are doing too many steps in one step

OpenStudy (anonymous):

what ?

OpenStudy (freckles):

I'm just saying because you did so many steps in one step you have made a couple of mistakes

OpenStudy (anonymous):

maybe..

OpenStudy (freckles):

no i'm sure

OpenStudy (freckles):

I think @Vocaloid was going to take you through the process

OpenStudy (anonymous):

Yeah idk

OpenStudy (freckles):

if i were you I would use the distributive property for the inner parts then move on to the outer parts

OpenStudy (dinamix):

@glitterythings [-2(3x-1)+3(5x-3^2)] find this first and after {7x-[-2(3x-1)+3(5x-3^2)]}

OpenStudy (anonymous):

I did that i thought , i might of skipped a step hold on

OpenStudy (anonymous):

Ill just figure it out myself thank you guys!

OpenStudy (freckles):

\[3-2{7x-[-2(3x-1)+3(5x-3^2)]} \\ \text{ like first things I would do } \\ -2(3x-1) \text{ and } 3(5x-3^2)\]

OpenStudy (anonymous):

I did that.. but i didnt put it in

OpenStudy (freckles):

that problem did not copy over to latex well

OpenStudy (dinamix):

[-2(3x-1)+3(5x-3^2)] = 9x-26

OpenStudy (anonymous):

oh i get what you guys are trying to tell me

OpenStudy (freckles):

\[3-2\{7x-[-2(3x-1)+3(5x-3^2)]\}\] there we go

OpenStudy (anonymous):

then it would be 3- 2{7x - [-2(3x-1) + 3(5x - 9)] right?

OpenStudy (freckles):

isn't that thing in the [ ] 9x-25?

OpenStudy (freckles):

@dinamix

OpenStudy (dinamix):

yup it is

OpenStudy (dinamix):

i make mistake

OpenStudy (dinamix):

{7x-[-2(3x-1)+3(5x-3^2)]} =7x-9x+25 =-2x+25

OpenStudy (freckles):

\[3-2\{7x-[-2(3x-1)+3(5x-3^2)]\} \\ 3-2\{7x-[-6x+2+15x-27]\} \\ 3-2\{7x-[9x-25]\} \\ \]

OpenStudy (freckles):

anyways @glitterythings maybe you can try from here?

OpenStudy (anonymous):

yeah i can !

OpenStudy (dinamix):

@freckles u find it 4x-(4!+4!-1)

OpenStudy (anonymous):

@dinamix im working on it rn :)

OpenStudy (anonymous):

i got 4x - 47

OpenStudy (dinamix):

yup @glitterythings this the answer ;D

OpenStudy (anonymous):

Ty :D

OpenStudy (anonymous):

i have another question it has to do the same thing , and i think i did it wrong lol.

OpenStudy (dinamix):

ok ask @glitterythings

OpenStudy (anonymous):

4y^2 - 3 {2y^2 -[(4y^2-y+2) - (3y^2 +2y-5)]}

OpenStudy (dinamix):

Start with this [(4y^2-y+2) - (3y^2 +2y-5)] u should try with me okey @glitterythings

OpenStudy (anonymous):

i am i solved, but answer is wrong. The answer is supposed to be y^2 - 9y +21

OpenStudy (anonymous):

ill re write again :)

OpenStudy (anonymous):

does everything become negative except for 5?

OpenStudy (dinamix):

u find it y^2-9y+21 ?

OpenStudy (anonymous):

not yet

OpenStudy (dinamix):

cuz i find like that

OpenStudy (anonymous):

you did? ugh idk what im doing wrong

OpenStudy (dinamix):

yup ,ok first i show my method start [(4y^2-y+2) - (3y^2 +2y-5)] this and tell me what u find okey

OpenStudy (dinamix):

i want see where is your mistake

OpenStudy (anonymous):

4y2 - 3y2 = 1 -y+2y = 1 2 + -5 = -3?

OpenStudy (anonymous):

i might be doing my math wrong

OpenStudy (dinamix):

yup i find your mistake u forget Distribution - on this equeation - (3y^2 +2y-5)

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

it would be -3y2 -2y +5?

OpenStudy (dinamix):

yup

OpenStudy (dinamix):

u have to focus @glitterythings (sorry if i was rude)

OpenStudy (anonymous):

so then [ y2 -3y + 7?]

OpenStudy (anonymous):

you arent.

OpenStudy (dinamix):

yup this

OpenStudy (dinamix):

{2y^2 -[(4y^2-y+2) - (3y^2 +2y-5)]} now second try for this

OpenStudy (dinamix):

@glitterythings

OpenStudy (anonymous):

oh i thought i did lol

OpenStudy (dinamix):

so u find it @glitterythings

OpenStudy (anonymous):

i think so.

OpenStudy (anonymous):

y2 + 3y -7?

OpenStudy (dinamix):

yup

OpenStudy (anonymous):

OMG I FOUND ITT!

OpenStudy (anonymous):

YAY THANKS SO MUCH!

OpenStudy (dinamix):

we didnt do it really u have to complete it -_-

OpenStudy (anonymous):

oh well i figured it out lol but okay lets complete it

OpenStudy (dinamix):

give me your steps okey

OpenStudy (anonymous):

so then you multiply by -3 4y2 -3 (y2+3y - 7) 4y2-3y2 - 9y +21 = y2 -9y+21

OpenStudy (anonymous):

thank you for being patient w me.

OpenStudy (dinamix):

yup this @glitterythings no problem u are welcome but i hope u understand the steps

OpenStudy (anonymous):

yeah i do, its like PEMDAS and you have to solve everything in brackets,

OpenStudy (dinamix):

ok that s make feel happy to help u ^^

OpenStudy (anonymous):

:)

OpenStudy (dinamix):

@glitterythings any other question ?